OpenStudy (anonymous):

Maria spins a penny 100 times and it lands head side up 62 times. Explain why Maria's experimental probability may be different from the theoretical probability of spinning a coin. @michele_laino

OpenStudy (michele_laino):

here we have to compute the experimental probability. Now in our experiment, we got 62 heads, so we have 62 favorable cases. Furthermore, we made 100 spins, so we have 100 possible cases, so, what is the experimental probability? \[\Large p = \frac{{favorable\;cases}}{{possible\;cases}} = ...?\]

OpenStudy (anonymous):

62 /100 = 0.62

OpenStudy (michele_laino):

that's right p=62%

OpenStudy (michele_laino):

now the theoretical probability is the same as in the previous exercise, so what is the theoretical probability?

OpenStudy (anonymous):

ummm idk

OpenStudy (michele_laino):

we have 2 faces (head and tail) so 2 possible cases, furthermore, we have 1 favorable case, namely the outcome head, so what is p?

OpenStudy (michele_laino):

\[\Large p = \frac{{favorable\;cases}}{{possible\;cases}} = ...?\]

OpenStudy (anonymous):

1/2

OpenStudy (michele_laino):

that's right! p=50%

OpenStudy (anonymous):

so the same like the other question

OpenStudy (michele_laino):

now the question is: the experimental probability is 62% whereas the theoretical probability is 50% why those probability are different each other, namely what is a possible cause?

OpenStudy (anonymous):

ummmmm

OpenStudy (michele_laino):

maybe that our coin is not a "honest" coin? Namely it is a rigged coin?

OpenStudy (anonymous):

maybe

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thanks!

OpenStudy (michele_laino):

Thanks! :)